# Testing Most Influential Sets

### v2

@@ 3.2 Extreme value distribution
    Relative-size sets: $k$ grows proportionally with $N,$ i.e., $k=pN$ for some $p\in(0,1).$
Both regimes appear in practical applications (see, e.g, Broderick et al., 2021; Kuschnig et al., 2021), but they yield fundamentally different asymptotic behavior with important implications for significance testing.
@@ 3.2.1 Constant-size Sets
Suppose $\mathbb{E}\left[X^{2}\right]<\infty,$ and that $X_{i},R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r}<\infty$. If $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ remains constant as $N\to\infty,$ then
@@
Let $C\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}R_{i}$ and $D\coloneqq\sum_{n=1}^{N}X_{n}^{2}.$ Notice that $C$ and $D^{-1}_{-{\mathbb{S}}}$ are asymptotically independent. Since $X_{i}$ and $R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r},$ their product satisfies $C\in\mathrm{MDA}(\Phi_{\xi}),$ with $\xi=\min\{\xi_{x},\xi_{r}\},$ and its upper tail behaves like the tail of $\max\{X_{i}R_{i}\}$ for $i\in{\mathbb{S}}_{k}^{\max}.$ Lemma 1 shows that the inverse sum $D^{-1}_{-{\mathbb{S}}}\in\mathrm{MDA}(\Lambda)$, and the product $CD^{-1}_{-{\mathbb{S}}}$ inherits the Fréchet behavior from $C$ by Lemma 2. See the Appendix for details. ∎
This result shows that for constant-size sets, $\Delta^{\max}$ exhibits heavy-tailed Fréchet behavior, implying that even a few observations can exert extreme influence with non-negligible probability.
@@
If the tail coefficients of both $X_{i}$ and $R_{i}$ is infinite, then
@@
3.2.2 Relative-size Sets
When the most influential set grows proportionally with the sample size, the central limit theorem dominates the asymptotic behavior:
Theorem 2 (EVD for relative-size sets).
If $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a central limit theorem (CLT) and $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ grows proportionally with $N,$ then
@@ 3.2.2 Relative-size Sets
When $\lvert{\mathbb{S}}_{k}^{\max}\rvert=pN,$ for $p\ll 1$, the numerator $C$ grows at the rate $\mathcal{O}(N).$ By the CLT, $C/\sqrt{N}\sim\mathcal{N}(\mu,\sigma^{2})$ as $N\to\infty$. Hence, the product $CD^{-1}_{-{\mathbb{S}}}$ lies in the maximum domain of attraction of the Gumbel distribution, following Lemma 2 and Corollary 2. See the Appendix for details. ∎
For relative-size sets, $\Delta^{\max}$ converges to a well-behaved Gumbel distribution with exponentially decaying tails, in contrast to the heavy-tailed Fréchet behavior of constant-size sets. This result holds regardless of the underlying distributions of $X$ and $R$ as long as the variance of $X_{i}\cdot R_{i}$ is finite.
3.3 Implementation and Computation
With theoretical results established, we turn towards practical implementation. Our procedure follows three steps:
@@ 3.3 Implementation and Computation
    Determine the relevant extreme value distribution. We select between the Gumbel and Fréchet families based on the hypothesized set size and the tail behavior of $X$ and $R$. We estimate tail coefficients using maximum likelihood estimation (MLE; Smith, 1985; Bücher & Segers, 2017). If $1/\max\{\xi_{x},\xi_{r}\}$ is sufficiently close to zero, we default to the Gumbel distribution (per Corollary 1 and Theorem 2). Otherwise, we use the Fréchet distribution with shape parameter $\xi=\max\{\xi_{x},\xi_{r}\},$ following Theorem 1.222Small values of $\xi$ correspond to extremely heavy-tailed distributions where the variance ($\xi\leqslant 2$) or even the mean ($\xi\leqslant 1$) become infinite. Such extreme cases pose practical challenges for statistical inference.
@@
    Estimate location and scale parameters. We estimate the location and scale parameters $a,b$ using the block maxima method (Coles, 2001; De Haan & Ferreira, 2006). We divide the sample (excluding ${\mathbb{S}}_{k}^{\max}$ for robustness) into $M$ blocks of size $N/M$, compute $\Delta^{\max}$ for each block, and use MLE based on these draws.
    Since selecting the maximum out of $N/M$ observations reduces the expected maximum compared to the full sample, in the Gumbel case a bias correction can be applied. More specifically, we know that
@@
    Perform hypothesis testing. We test the null hypothesis $H_{0}$ that the observed influence reflects natural sampling variation against the $H_{1}$ of excessive influence. Using the estimated parameters, we compute the $p$-value as $P(\Delta^{\max}\geqslant\delta_{\text{obs}})$ where $\delta_{\text{obs}}$ is the observed maximum influence.
Computational Efficiency.
Our procedure is computationally convenient, allowing for application to large and varied datasets. The maximum likelihood steps are simple and well-behaved, optimizing over only two parameters in the Gumbel case. The primary computational constraint stems from finding most influential sets — we need to approximate $\Delta^{\max}$ for the $M$ block maxima estimates (Price et al., 2022). For computational tractability, we use an adaptive greedy algorithm (Hu et al., 2024; Kuschnig et al., 2021) with complexity $\mathcal{O}(Mk)$ and considerably reduced runtime from our closed-form influence formula for sets in Proposition 1.
@@ 4 Experiments
We validate our theoretical predictions and demonstrate practical utility through controlled simulations and real-world applications spanning economics, biology, and machine learning.
@@ 4.1 Simulation Study
Figure 1: Illustration of our methodology on a simple linear regression with a moderately influential observation. Panel A depicts observations, estimated regression lines with and without the influential point, and conditional significance regions at the 10, 5, and 1% levels (dotted lines). Panel B shows the extreme value analysis: a histogram of block maxima, fitted Gumbel distribution with (solid) and without (dashed) bias correction, and the resulting $p$-value for the observation of interest.
@@ 4.1.1 Convergence to Extreme Value Distributions
We verify that maximal influence converges to the predicted extreme value distributions. Across for scenarios combining standard Normal and $t(5)$ distributions for $X,R,$ we simulate $1,000$ datasets of sizes $N\in\{100,200,500,1000,2000\}.$
Table 1 shows inverse shape estimates $\hat{\xi}^{-1}$ for the setup with $N=100.$ As predicted by theory, the Normal-Normal case yields clear Gumbel behavior ($\xi^{-1}=0$) for all sample sizes (see Table A1). The heavy-tailed cases exhibit predicted Fréchet behavior, with shape estimates matching theoretical predictions ($\xi^{-1}=0.2$), confirming the the finite sample applicability of Theorem 1. Particularly the convergence of the t(5)t(5)–Normal case has slower convergence rates to the limiting distribution due to the instability of $D^{-1}_{\mathbb{S}}$ for small samples (see Figure A2).
Table 1: Inverse Shape Estimates for Different Distributions
Mean Std.Dev. Q25 Median Q75
Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
$1,000$ repetitions for $N=100$ observations.
4.1.2 Location and Scale Estimation
We evaluate whether estimation of location and scale parameters with block maximum MLE can accurately capture the true distribution. (Results are provided in Figure A3.) The finite sample adjusted location parameter works well and shows convergence to the true value, while the scale exhibit slight downward bias that is consistent with known limitations of MLE for EVDs (Dombry & Ferreira, 2019). Panel (a) in Figure A3 summarises the effectiveness of the MLE in the Gumbel case, allowing us to conduct hypothesis tests.
@@ 4.2 Applications
We investigate several real-world datasets — two applications from economics and biology, as well as four well-known machine learning benchmarks.
4.2.1 Economic Development and Geography
We re-examine the controversial finding that rugged terrain benefits African economies when compared to the rest of the world (Nunn, 2020). Kuschnig et al. (2021) identify the Seychelles, coupled with any of Rwanda, Lesotho, Eswatini, and the Comoros, as influential, removing significance of the estimate of interest.
Our results decisively resolve this controversy. Table 2 reveals the Seychelles as excessively influential on $\hat{\theta}_{\text{rugged}}$, both individually ($p<0.001$) and in combination with other outliers except for Lesotho. This confirms the suspected confounding from the size of nations, lending statistical rigor to prior concerns.
@@ 4.2.1 Economic Development and Geography
4.2.2 Sparrow Morphology — Head and Beak Size
We analyze the relation between head and tarsus length in saltmarsh sparrows, based on measurements of $N=1295$ sparrows with known outliers (Gjerdrum et al., 2008; Zuur et al., 2010). The baseline regression yields $\hat{\theta}=0.011$ with a standard error of $(.030),$ implying a relation that is statistically indistinguishable from zero.
However, a curious data point moves the estimate to $0.219(.029)$, turning the estimate significantly positive. An additional data point further moves the estimate to $0.288(.032).$ These extreme impacts from a vanishing fraction of the sample are deemed excessive by our approach at any conventional significance level (both p<0.001p<0.001). [^3] A possible explanation for this excessive influence are data entry errors: The first observation (an outlier in both head and tarsus size) may have the two (adjacent) features mixed up — when swapped, they would fit well into overall averages. The second observation (an outlier in one feature) stands out with both values being equal up to the one significant digit.
4.2.3 Machine Learning Benchmarks
@@
  • Law School ($N=20,800$): We examine the coefficient for the ‘Other’ race indicator, with $378$ relevant samples. We consider two sets: $77$ data points that move the estimate from 0.0412(.0144)-0.0412\ (.0144) to 0.1117(.0159),0.1117\ (.0159), creating a significant estimate with flipped sign, and $17$ data points that reduce the estimate to 0.0223(.0097).-0.0223\ (.0097). Our approach indicates that the influence larger set’s influence falls within expected variation, while the smaller set exhibits statistically excessive influence ($p=0.019$).
@@
  • Boston Housing ($N=506$): We focus on the effect of crime rate on house values. The baseline (highly significant) coefficient 0.1080(.0329)-0.1080\ (.0329) is rendered insignificant at 0.0352(.0556)-0.0352\ (.0556) after excluding just $6$ observations. In this case, the underlying EVD is Frechét with inverse shape $\xi^{-1}=0.29$ due to the heavy tail of the crime variable. The set’s influence is highly significant ($p=0.001$), indicating excessive influence.
  • Communities & Crime ($N=1,994$): We investigate $2$ and $2$ data points with substantial influence on the relation between race and crime rates. The complete set is not extreme, as the points cancel each other out. After exclusion, the first subset of two increases the coefficient by more than 22%, which is deemed excessive $p<0.001.$ When re-estimating after their exclusion, the second set decreases the estimate by more than 10% and is deemed excessive at the 5% level ($p=0.014.$ (See Table A2 for details.)
@@ 5 Discussion
Our analysis provides the first rigorous statistical framework for assessing when most influential sets represent genuine problems rather than natural sampling variation. By establishing that maximal influence follows predictable extreme value distributions, we enable practitioners to move beyond ad-hoc rules and domain-specific judgment. The key insight is that maximal influence fundamentally depends on the nature of the sets considered and the tail behavior of the underlying data.
This work addresses a critical gap in interpretable machine learning, where theoretical foundations for influential set analysis have been lacking. We show that maximal influence depends on the tail properties of the underlying data—when tails are heavy, influence patterns become unpredictable, requiring robust estimation methods or acceptance of the inherent instability. For distributions with moderate tail behavior, our results provide tight theoretical bounds.
Our approach integrates naturally with the literature on influential set selection (Broderick et al., 2021; Fisher et al., 2023; Hu et al., 2024), which has lacked conclusive theoretical guidance. While influential sets share connections with established diagnostics such as Cook’s distance and leverage scores, or methods such as robust regression (Huber & Ronchetti, 2009), they fill a unique role. By directly relating to quantities of interest, they allow practitioners to discover and analyze sets of data that deviate from dominant patterns and yield insights that other methods obscure (see, e.g., Kuschnig et al., 2021, for comparisons). Our results provide the theoretical foundations for analyzing these influential sets. [^4] We can also clarify the applicability of the common $2\sqrt{N}$ threshold for coefficient influence (Belsley et al., 1980). While imprecise for most influential observations — missing the selection procedure that necessitates extreme value theory — it proves asymptotically accurate for randomly selected observations.
5.1 Limitations and Future Work
Our analysis operates within linear regression — a foundational setting for theory and modern ML, but limited to contexts where interpretability is valued (Rudin, 2019; Roscher et al., 2020). While we focus on regression coefficients, an extension predictions follows trivially through $\hat{Y}=X\hat{\beta}$. Extending our results to generalized linear, tree-based, or non-parametric models, requires further developments.
The asymptotic arguments presented here provide important insights, but come with inherent limitations. Our analysis leverages independence between features and the residuals, which can be restrictive in practice where dependence affects influence patterns. While we address these concerns in controlled settings, the gap between theory and finite-sample behavior warrants investigation.
Several methodological improvements could strengthen practical performance. Estimation of extreme value parameters could be enhanced through setting-specific methods and improved bias correction (Dombry & Ferreira, 2019). The efficient selection of most influential sets themselves remains an open challenge (Hu et al., 2024; Huang et al., 2025) with direct implications for our procedure.
5.2 Broader Implications
@@
This work enables more reliable and transparent decision-making across domains where linear models remain the method of choice. Principled tools for understanding data points that drive model behavior are crucial for building trustworthy systems. Applications span fairness assessments, where influential subsets can reveal algorithmic bias, to causal inference settings, including randomized controlled trials quasi-experimental econometric analyses where small data subset can fundamentally alter estimates.
@@
Importantly, we reframe influence as a natural and informative feature of data that requires appropriate treatment rather than a probem to be fixed. Influential sets can represent genuine heterogeneity or important edge cases that should inform model development. This perspective emables more nuanced approaches to data processing and model validation, where more information is preserved and assessed through principled statistical inference rather than discarded based on rules of thumb.

### v3

@@ 3.1 Extreme Value Distributions
    Growing-size sets: $k$ grows as $N\to\infty$ and $k/N\to 0$
Both regimes have been considered in practical applications (see, e.g, Broderick et al., 2023; Kuschnig et al., 2021), and — as we will show next — they yield fundamentally different asymptotic behavior with important implications for the interpretation of influence.
@@ 3.1.1 Constant-size Sets
Suppose $\mathbb{E}\left[X^{2}\right]<\infty,$ and that the heavier tail of $X_{i},R_{i}$ decays at polynomial speed with coefficients $\xi_{x},\xi_{r}$ where $\min\{\xi_{x},\xi_{r}\}<\infty.$ If $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ remains constant as $N\to\infty,$ then
@@
Let $C\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}R_{i}$ and $D\coloneqq\sum_{n=1}^{N}X_{n}^{2}.$ Notice that $C$ and $D^{-1}_{-{\mathbb{S}}}$ are asymptotically independent. Since $X_{i}$ and $R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r},$ their product $C\in\mathrm{MDA}(\Phi_{\xi}),$ with $\xi=\min\{\xi_{x},\xi_{r}\},$ since its upper tail behaves like the tail of $\max\{X_{i}R_{i}\}$ for $i\in{\mathbb{S}}_{k}^{\max}.$ Lemma 1 shows that the inverse sum $D^{-1}_{-{\mathbb{S}}}\in\mathrm{MDA}(\Lambda)$, and the product $CD^{-1}_{-{\mathbb{S}}}$ inherits the Fréchet behavior from $C$ by Lemma 2. ∎
This result shows that, for constant-size sets, $\Delta^{\max}$ inherits tail behavior from the heavier tail of $R$ and $X.$ If one of them is sufficiently heavy-tailed, even small sets can exert extreme influence with non-negligible probability. Corollary 1 simplifies Theorem 1 in absence of heavy tails.
@@
If the tail coefficients of both $X_{i}$ and $R_{i}$ are infinite, then
@@
3.2 Growing Sets
When the most influential set grows ‘slow enough’ with the sample size, the central limit theorem (CLT) dominates the asymptotic behavior:
Theorem 2 (EVD for growing sets).
If $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a CLT and $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ grows at $o(N)$ but faster than $O(1),$ then
@@ 3.2 Growing Sets
Let $m_{k}=\lvert{\mathbb{S}}_{k}^{\max}\rvert$. By assumption, $m_{k}\to\infty$ and $m_{k}=o(N)$ as $N\to\infty$. The numerator, $C,$ can be written as a partial sum of $m_{k}$ terms and grows at rate $\mathcal{O}(m_{k})$. Since $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a CLT, we obtain $\left(C-\mathbb{E}[C]\right)/\sqrt{m_{k}}\;\overset{d}{\longrightarrow}\;\mathcal{N}(\mu,\sigma^{2})$ as $N\to\infty$. Hence, following Lemma 2 and Corollary 2, the product $CD^{-1}_{-{\mathbb{S}}}$ lies in the maximum domain of attraction of the Gumbel distribution. ∎
This reveals a fundamental distinction: constant-size sets are dominated by the heaviest tail, while, for growing sets, $\Delta^{\max}$ converges to a well-behaved Gumbel distribution with exponentially decaying tails. This result holds regardless of the underlying distributions of $X$ and $R$ as long as the variance of $X_{i}\cdot R_{i}$ is finite.
3.3 Implementation
With the theoretical results established, we turn towards practical implementation. Assuming there is a most influential set of interest, [^3] This set can be obtained with any of the methods in the literature (Broderick et al., 2023; Kuschnig et al., 2021; Freund and Hopkins, 2023), and could, e.g., be the smallest set that achieves a sign-flip — a commonly considered heuristic cutoff. If no such heuristic is used and multiple tests are considered (e.g., for different coefficients or sets sizes), a multiple testing correction should be applied. Note that EVT controls the implicit search for the most influential set over the $\binom{N}{k}$ possible subsets. our procedure follows three steps:
@@ 3.3 Implementation
    Choose the EVD family. Our theoretical results guide the decision between the Gumbel and Fréchet families, which based on the hypothesized set size and the tail behavior of $X$ and $R$. For the latter, we can estimate tail coefficients using maximum likelihood estimation (MLE; Smith, 1985; Bücher and Segers, 2017). If $1/\xi$ is sufficiently close to zero, we default to the Gumbel distribution (per Corollary 1 and Theorem 2). Otherwise, we use the Fréchet distribution with shape parameter $\xi,$ following Theorem 1.444Small values of $\xi$ correspond to extremely heavy-tailed distributions where the variance ($\xi\leqslant 2$) or even the mean ($\xi\leqslant 1$) become infinite. Such cases pose practical challenges for statistical inference, and imply that arbitrarily large influence is possible in our case.
@@
    Estimate EVD parameters. With the EVD known, we estimate its location and scale parameters $a,b$ e.g., using the block maxima method (Coles, 2002; de Haan and Ferreira, 2006). For this, we divide the sample (excluding ${\mathbb{S}}_{k}^{\max}$ for robustness) into $M$ blocks of size $N/M$, compute $\Delta^{\max}$ for each block, and use MLE based on these draws. Since selecting the maximum out of $N/M$ observations reduces the expected maximum compared to the full sample, a bias correction can be applied for the Gumbel distribution, leveraging that the densities for sizes $N$ and $N/M$ are related by
@@
    Perform hypothesis test. Finally, we test the null hypothesis $H_{0}$ that the observed influence reflects natural sampling variation against the alternative $H_{1}$ of excessive influence. Based on the estimated parameters, we can simply compute the $p$-value as $P(\Delta^{\max}\geqslant\delta_{\text{obs}})$ where $\delta_{\text{obs}}$ is the observed maximum influence.
Computation.
Thanks to Proposition 1, our procedure is practical and computationally convenient, allowing for application to large and varied datasets. While accurately estimating Extreme Value Distributions (EVDs) is a recognized and separate challenge in its own right, the maximum likelihood steps within our framework are simple and well-behaved, optimizing over only two parameters in the Gumbel case. The primary computational constraint instead stems from finding most influential sets — we need to approximate $\Delta^{\max}$ for the $M$ block maxima estimates. For computational tractability, we use an adaptive greedy algorithm (Hu et al., 2024; Kuschnig et al., 2021) with complexity $\mathcal{O}(Mk)$ and an efficient implementation based on our closed-form influence formula for sets.
@@ 4 Experiments
In this section, we validate our theoretical predictions, investigate convergence in small samples, and demonstrate their practical relevance and utility through real-world applications spanning economics, biology, and machine learning.
@@ 4.1 Simulation Study
We begin with a controlled setting, where we (i) illustrate, (ii) investigate convergence for small samples, and (iii) evaluate empirical estimation.
Illustration.
@@
Figure 2: Illustration of our methodology on a simple linear regression with a moderately influential observation. Panel A depicts observations, estimated regression lines with and without the influential point, and conditional significance regions at the 10, 5, and 1% levels (dotted lines). Panel B illustrates the extreme value analysis: a histogram of block maxima in the background, fitted Gumbel distributions with (solid) and without (dashed) bias correction, and the resulting $p$-value for the observation of interest.
@@ 4.1.1 Convergence to Extreme Value Distributions
Next, we verify that maximal influence converges to the predicted extreme value distributions. We consider four scenarios based on combinations of the standard Normal and $t(5)$ distributions for $X,R.$ For each scenario, we simulate $1,000$ datasets of sizes between $N=20$ and $N=1000,$ and compare the empirical parameter estimates with the theoretical prediction. Overall, we find rapid convergence, implying that our theoretical predictions are applicable with small samples.
Figure 3 shows convergence of the scenarios to the predictions, which are indicated by dashed horizontal lines. (Details are provided in Table A1 of the Appendix.) All scenarios reliably converge for moderate sample sizes. The Normal-Normal scenario is insignificantly different from Gumbel behavior ($\xi^{-1}=0$) for $N\geqslant 50,$ and the heavy-tailed cases also exhibit the predicted Fréchet behavior ($\xi^{-1}=0.2$). Notably, the t(5)t(5)–Normal case converges at slower rates, likely due to the relative instability of the inverse $\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}_{\mathbb{S}}$ in small samples. Overall, the simulation results support the applicability of Theorem 1 in small samples.
Figure 3: Convergence of empirical samples to the limiting distribution with increasing sample sizes ($N\in\{20,30,50,75,100,150,200,300,500,1000\},$ note the non-linear scale) for four different scenarios. Dots indicate estimated means from $1000$ repetitions, while the shaded area indicates $\pm 1\mathrm{SD}$. The Normal–Normal and the $t(5)t(5)$ cases quickly converge to the theoretical predictions of $\xi^{-1}=0$ and $\xi^{-1}=0.2,$ while the mixed scenarios converge marginally slower.
Location and Scale Estimation.
Next, we evaluate whether block maximum MLE accurately captures the location and scale parameters for empirical testing. Results are provided in Figure A1 of the Appendix. We find that bias-corrected estimation of the location parameter works well, while the scale estimate is consistent but exhibits a minor downward bias that disappears asymptotically (consistent with known limitations of the MLE; see Dombry and Ferreira, 2019). For our goal of hypothesis testing, the overall distribution and its quantiles are recovered effectively.
@@ 4.2 Applications
We investigate several real-world datasets — two applications from economics and biology, and four machine learning benchmarks — and provide the first conclusive investigation of influence.
Economic Development and Geography.
We re-examine the controversial finding that rugged terrain benefits African economies when compared to the rest of the world (Nunn and Puga, 2012). Kuschnig et al. (2021) call the influence of the Seychelles, which remove significance of the estimate of interest when coupled with any of Rwanda, Lesotho, Eswatini, and the Comoros. However, lacking the statistical framework that we provide, they were not able to test whether this level of influence should be deemed excessive or not.
We can now decisively resolve this controversy. Table 1 reveals the Seychelles as excessively influential on $\hat{\theta}_{\text{rugged}}$, both individually ($p<0.001$) and in combination with other outliers, except for Lesotho. This substantiates the suspected confounding from the size of nations (Kuschnig et al., 2021), lending statistical rigor to prior concerns and calling into question the differential relation between ruggedness and income in Africa.
@@
Sparrow Morphology — Big Heads and Beaks.
We analyze the relation between head and tarsus length in saltmarsh sparrows, based on measurements of $N=1,295$ sparrows with known outliers (Gjerdrum et al., 2008; Zuur et al., 2010). The baseline regression yields $\hat{\theta}=0.011$ with a standard error of $(.030),$ implying a relation that is statistically indistinguishable from zero.
However, a curious data point moves the estimate to $0.219(.029)$, turning the estimate significantly positive. An additional data point further moves the estimate to $0.288(.032).$ These extreme impacts from a vanishing fraction of the sample are deemed excessive by our approach at any conventional significance level (both p<0.001p<0.001). [^5] One possible explanation for this excessive influence are data entry errors: The first observation (an outlier in both head and tarsus size) may have the two (adjacent) features mixed up — when swapped, they would fit well into overall averages. The second observation (an outlier in one feature) stands out with both values being equal up to the one significant digit.
Machine Learning Benchmarks.
@@
  • Law School ($N=20,800$): We examine the coefficient for the ‘Other’ race indicator, with $378$ relevant samples. We consider two sets: $77$ data points that move the estimate from 0.0412(.0144)-0.0412\ (.0144) to 0.1117(.0159),0.1117\ (.0159), creating a significant estimate with flipped sign, and $17$ data points that reduce the estimate to 0.0223(.0097).-0.0223\ (.0097). Our approach indicates that the larger set’s influence falls within expected variation, while the smaller set exhibits statistically excessive influence ($p=0.019$).
@@
  • Boston Housing ($N=506$): We focus on the effect of crime rate on house values. The baseline (highly significant) coefficient 0.1080(.0329)-0.1080\ (.0329) is rendered insignificant at 0.0352(.0556)-0.0352\ (.0556) after excluding just $6$ observations. In this case, the underlying EVD is Fréchet with inverse shape $\xi^{-1}=0.29$ due to the heavy tail of the crime variable. The set’s influence is highly significant ($p=0.001$), indicating excessive influence.
  • Communities & Crime ($N=1,994$): We investigate $2$ and $2$ data points with substantial influence on the relation between race and crime rates. The complete set is not extreme, as the points cancel each other out. After exclusion, the first subset of two increases the coefficient by more than 22%, which is deemed excessive $p<0.001.$ When re-estimating after their exclusion, the second set decreases the estimate by more than 10% and is deemed excessive at the 5% level ($p=0.014$). (See Table A2 for details.)
@@ 5 Discussion
We develop the first rigorous statistical framework for assessing when most influential sets represent genuine problems rather than natural sampling variation. By deriving the extreme value distribution of maximal influence, we allow practitioners to replace ad-hoc sensitivity checks with principled statistical decision rules.
Our key insight is how maximal influence fundamentally depends on set size and tail behavior. For constant-size sets with polynomial tails, maximal influence follows a heavy-tailed Fréchet distribution, implying that extreme influence can be arbitrarily large. For growing sets or exponential-tailed data, maximal influence converges to a well-behaved Gumbel distribution.
These results address a critical gap in interpretable machine learning. While recent work has developed methods to identify influential sets (Broderick et al., 2023; Freund and Hopkins, 2023; Hu et al., 2024), no formal theory existed to determine when their influence is excessive. Our framework provides the long-missing theoretical foundation that enables rigorous statistical inference for influential observations and sets (first discussed by Cook, 1979).
We can clarify the applicability of heuristics that are commonly used for identifying excessive influence, and provide conclusive answers when excessively influential sets are suspected (such as the Seychelles in the ruggedness example; see Kuschnig et al., 2021). The $2/\sqrt{N}$ threshold for coefficient influence (Belsley et al., 1980), e.g., is asymptotically accurate for randomly selected observations, but is too restrictive for most influential observations, where the selection procedure necessitates extreme value theory.
Practical Recommendations.
In general, most influential sets hold valuable information for inference. Our test is deliberately conservative, controlling Type I errors (false claims of excessive influence) at the cost of some Type II errors (failing to detect truly excessive influence). This reflects our view that influential sets are a natural feature of data and not a problem to be eliminated. When our test identifies an excessively influential set, however, we recommend:
  1. 1.
    Investigate mechanism. Document the set and investigate why it differs — it may convey genuine heterogeneity, data quality issues, unobserved confounding, or important edge cases not addressed by the model.
  2. 2.
    Handle appropriately. We argue that an excessively influential set warrants separate analysis; exclusion can be considered if it reflects measurement error or outliers that are irrelevant for the pattern of interest.
  3. 3.
    Report transparently. State the set, decision, and test outcome; if conclusions hinge on the set, report both and discuss why. We advise against trimming or winsorizing to force alignment with the remaining data; these transformations create artificial data that may obscure and distort rather than illuminate underlying relations.
Limitations and Future Work.
Our analysis focuses on linear regression — foundational for theory and modern ML methods, but limited to contexts where interpretability is valued (Rudin, 2019). Extensions to generalized linear models, tree-based methods, or non-parametric estimators require further developments. Our asymptotic arguments leverage independence between features and residuals, which may be restrictive when dependence affects influence patterns. Explicitly addressing dependence between influence components through the selected set and across sets is possible through generalizations. While simulations show that small-sample behavior quickly converges to asymptotic predictions (even at $N=100$), further investigation of the theory-practice gap is warranted.
Practicality of our approach hinges on estimating the extreme value null, and, hence, approximating maximum influence. Finite-sample estimation of tail behavior and EVD parameters can be delicate, and improving these estimators would directly sharpen $p$-values. Two options include leveraging domain-specific information and improved bias correction methods (Dombry and Ferreira, 2019; Oorschot and Zhou, 2020). Advances in finding influential sets remain an active research area (Hu et al., 2024; Huang et al., 2025), and could substantially improve block-maxima, reduce runtime, and broaden applicability.
Broader Implications.
@@
Our framework enables more reliable decision-making across domains where linear models remain the method of choice. Principled tools for understanding data points that drive model behavior are crucial for building trustworthy systems. Applications span fairness assessments — where influential sets can reveal algorithmic bias — to causal inference settings, such as randomized controlled trials or quasi-experimental econometric analyses where small data subsets can fundamentally alter estimates.
@@
Importantly, we reframe influence as a natural feature of data requiring appropriate treatment rather than a problem to be fixed. Influential sets can represent genuine heterogeneity or important edge cases that should inform model development. This perspective enables more nuanced approaches to data analysis, where information is preserved and assessed through principled statistical inference rather than discarded based on rules of thumb.